A Level Mathematics 9709 Paper 53, May/June 2023

Question paper and mark scheme50 marks7 questions

Solve it on ExamPal

Practise

Untimed. Write on the paper and check as you go.

Start

Exam

Timed, with a clock you set, like sitting the real paper.

Start

Opens full screen.

Papers open in

Opening the question paper

Read as textThe questions and the mark scheme as plain text. Easier to read on a phone.

Question paper

9709/53 · May/June 2023

Download PDF
1

Two fair coins are thrown at the same time repeatedly until a pair of heads is obtained. The number of throws taken is denoted by the random variable X.

a

State the value of E(X).

[1]
b

Find the probability that exactly 5 throws are required to obtain a pair of heads.

[1]
c

Find the probability that fewer than 7 throws are required to obtain a pair of heads.

[2]
2

Anil is a candidate in an election. He received 40% of the votes. A random sample of 120 voters is chosen. Use an approximation to find the probability that, of the 120 voters, between 36 and 54 inclusive voted for Anil.

[5]
3

The random variable X takes the values 1, 2, 3, 4. It is given that P(X=x)=kx(x+a), where k and a are constants.

a

Given that P(X=4)=3P(X=2), find the value of a and the value of k.

[4]
b

Draw up the probability distribution table for X, giving the probabilities as numerical fractions.

[1]
c

Given that E(X)=3.2, find Var⁡(X).

[2]
4

The times taken, in minutes, to complete a cycle race by 19 cyclists from each of two clubs, the Cheetahs and the Panthers, are represented in the following back-to-back stem-and-leaf diagram.

CheetahsPanthers
9874
87320868
987917899
6533110234456
98211128
41206

Key: 7∣9∣1 means 97 minutes for Cheetahs and 91 minutes for Panthers

a

Find the median and the interquartile range of the times of the Cheetahs.

[3]
b

The median and interquartile range for the Panthers are 103 minutes and 14 minutes respectively. Make two comparisons between the times taken by the Cheetahs and the times taken by the Panthers.

[2]
c

Another cyclist, Kenny, from the Cheetahs also took part in the race. The mean time taken by the 20 cyclists from the Cheetahs was 99 minutes. Find the time taken by Kenny to complete the race.

[3]
5

Jasmine throws two ordinary fair 6-sided dice at the same time and notes the numbers on the uppermost faces. The events A and B are defined as follows. A: The sum of the two numbers is less than 6. B: The difference between the two numbers is at most 2.

a

Determine whether or not the events A and B are independent.

[4]
b

Find P(B∣A′).

[3]
6

The mass of grapes sold per day by a large shop can be modelled by a normal distribution with mean 28 kg. On 10% of days less than 16 kg of grapes are sold.

a

Find the standard deviation of the mass of grapes sold per day.

[3]
b

The mass of grapes sold on any day is independent of the mass sold on any other day. 12 days are chosen at random. Find the probability that less than 16 kg of grapes are sold on more than 2 of these 12 days.

[3]
c

In a random sample of 365 days, on how many days would you expect the mass of grapes sold to be within 1.3 standard deviations of the mean?

[4]
7
a

Find the number of different arrangements of the 10 letters in the word CASABLANCA in which the two Cs are not together.

[3]
b

Find the number of different arrangements of the 10 letters in the word CASABLANCA which have an A at the beginning, an A at the end and exactly 3 letters between the 2 Cs.

[3]
c

Five letters are selected from the 10 letters in the word CASABLANCA. Find the number of different selections in which the five letters include at least two As and at most one C.

[3]

Mark scheme

9709/53 · May/June 2023

Download PDF

1a · 1 mark

  • B1

    [P(HH)=14][E(X)=]4

1b · 1 mark

  • B1

    [P(X=5)=(34)4(14)=]0.0791

1c · 2 marks

  • M1

    [P(X<7)=]1−(34)6 or 14+34×14+3242×14+…+3545×14

  • A1

    =33674096,0.822

2 · 5 marks

  • B1

    Mean=120×0.4=48 Var=120×0.4×0.6=28.8

  • M1

    Substituting their μ and σ into one ±standardisation formula (any number for 35.5 or 54.5), condone σ2 and σ.

  • M1

    Using continuity correction 35.5, 36.5 or 53.5, 54.5 once in their standardisation formula.

  • M1

    [=P(−2.3292<Z<1.211)=]0.8871+0.9900−1

  • A1

    =0.877

3a · 4 marks

  • M1

    [P(X=4)=3P(X=2)] 4k(4+a)=3×2k(2+a) 16k+4ak=12k+6ak

  • A1

    a=2

  • M1

    3k+8k+15k+24k=1

  • A1

    k=150

3b · 1 mark

  • B1
    X1234
    P(X)350,0.06850,0.161550,0.32450,0.48

3c · 2 marks

  • M1

    Var(X)=350×12+850×22+1550×32+2450×42−3.22

  • A1

    [=11.08−3.22=]0.84[0],2125

4a · 3 marks

  • B1

    Median = 99 [minutes]

  • M1

    [IQR=]106−83

  • A1

    23 [minutes]

4b · 2 marks

  • B1

    The times for the Cheetahs are faster than the times for the Panthers

  • B1

    The times for the Cheetahs are more spread than the times for the Panthers

4c · 3 marks

  • B1

    [Total time including Kenny = 99×20= ]1980

  • M1

    [Kenny's time =] 1980−1862

  • A1

    = 118 [minutes]

5a · 4 marks

  • B1

    P(A)=1036 P(B)=2436

  • B1

    P(A∩B)=836

  • M1

    1036×2436

  • A1

    =527,0.185[≠836] Events are not independent

5b · 3 marks

  • M1

    [P(B∩A′)=1636,0.4444] or their P(B) – their P(A∩B) seen as numerator or denominator of conditional probability fraction.

  • M1

    [P(A′)=1−1036]=2636,0.7222 or 1− their P(A) seen as denominator of conditional probability fraction.

  • A1

    =813

6a · 3 marks

  • B1

    ±1.282 seen, cao – critical value.

  • M1

    Use of the ±standardisation formula with 16, 28, σ and a z-value (not 0.1, 0.9, 0.282, 0.5398, 0.8159) equated to a z-value.

  • A1

    σ=9.36

6b · 3 marks

  • M1

    One term 12Cx(p)x(1−p)12−x,0<p<1.x≠0,1,2.

  • A1

    Correct expression, accept unsimplified, no terms omitted leading to final answer.

  • B1

    0.111

6c · 4 marks

  • B1

    Identifying at least one of −1.3 or 1.3 as the appropriate z-values.

  • M1

    Calculating the appropriate probability area from 2 symmetrical z-values (leading to their final answer, expect > 0.5).

  • A1

    =0.806,504625

  • B1

    [In 365 days 0.8064×365] =294 or 295

7a · 3 marks

  • M1

    10!a!b!−c,a≠b,a=1,2,b=1,4, with c being a positive integer.

  • M1

    d−e!4!,e=8,9,10, with d being a positive integer.

  • A1

    = 60480

7b · 3 marks

  • M1

    6!2!×s, with s being a positive integer.

  • M1

    t!r!×4,r=1,2,3 and t=8,7,6.

  • A1

    1440

7c · 3 marks

  • B1

    Correct number of ways for identified scenarios of 2 or 3 As, accept unsimplified, www.

  • M1

    Add 3 values for 2, 3 and 4 As, no additional, incorrect or repeated scenarios. Accept unsimplified.

  • A1

    25

Questions

  1. 1
    Question 14 marks
    Mark scheme ›
  2. 2
    Question 25 marks
    Mark scheme ›
  3. 3
    Question 37 marks
    Mark scheme ›
  4. 4
    Question 48 marks
    Mark scheme ›
  5. 5
    Question 57 marks
    Mark scheme ›
  6. 6
    Question 610 marks
    Mark scheme ›
  7. 7
    Question 79 marks
    Mark scheme ›

More 9709 papers

Other papers from May/June 2023

Solve it on ExamPalStart